# Disproving the quasi-uniformity of the Halton sequences and some of Halton-type sequences

### v1

@@ Abstract.
In this short article, we prove that the Halton sequence, one of the most well-known low-discrepancy sequences, is not quasi-uniform in any dimension $d\geq 2$ with any pairwise relatively prime bases. We further disprove the quasi-uniformity of some Halton-type sequences, including the Faure sequence, which provides an alternative proof of the known results.
@@ 1. Introduction
The Halton sequence is evidently one of the most well-known low-discrepancy sequences in the unit cube $0,1]^{d}$ for any dimension $d\geq 1$ [LP14, Nie92]. Throughout this paper, for an integer $b\geq 2$, we abbreviate the base-$b$ representation of $n=n_{0}+n_{1}b+n_{2}b^{2}+\cdots\in\mathbb{N}_{0}$ as $(n_{0},n_{1},n_{2},\ldots)_{b}$, where $\mathbb{N}_{0}$ denotes the set of non-negative integers. The radical inverse function $\varphi_{b}$ in base $b$ is then defined by
@@
The discrepancy of the Halton sequence was studied already by Halton himself in [Hal60], who proved an upper bound of $\mathcal{O}(N^{-1}(\log N)^{d})$ for the star discrepancy of the initial $N$ points of the sequence. Much later, Atanassov [Ata04] provided a refined analysis, which led to a substantial improvement of the implied constant in this $\mathcal{O}$-bound. While Halton’s original constant grows super-exponentially fast in the dimension $d$, Atanassov’s constant decreases to 0 super-exponentially. In [Le16a, Le16b], Levin proved a matching-order lower bound for the star discrepancy of the Halton sequence. Moreover, the tractability property of the weighted star discrepancy of Halton sequences has also been studied in [HPT19].
With these results, together with the classical Koksma–Hlawka inequality [KN74, Chapter 2], the use of Halton sequences in high-dimensional numerical integration is theoretically well justified. However, Halton sequences have also been employed as sampling nodes for scattered data approximation—particularly in the context of radial basis function approximation (see, for instance, [FM12, MS17, SSW23]). In such applications, not the low discrepancy but rather the geometric properties of the point set, such as the coverage of the underlying domain and the spacing between points, are crucial, as they contribute to bounds on the approximation error and to numerical stability [DS10, S95, W05, WSH21]. These geometric properties can be quantified in terms of the covering radius and the separation radius of the point set, naturally leading to the notion of quasi-uniformity.
@@
see, for instance, [DGS25, Remarks 2.4 & 2.5]. Thus, the quasi-uniformity requires that both radii decay at the optimal rate of order $N^{-1/d}$. As shown, for instance, in [Nie92, Theorem 6.12], it is known that for $\mathcal{S}$ being the Halton sequence in any relatively prime bases $b_{1},\ldots,b_{d}\geq 2$,
@@
In passing, it is known that the first point $\boldsymbol{x}_{0}$, which lies at the origin, is well separated from the subsequent points of the Halton sequence [HKZ05, Owe06]. In the present context, this result can be rephrased as the existence of a constant $c_{d}>0$, depending only on the bases $b_{1},\ldots,b_{d}$, such that
@@
One of the main purposes of this article is to disprove the quasi-uniformity of the Halton sequences in any dimension $d\geq 2$. This is done in Section 2 by proving that the separation radius of the Halton sequence indeed decays faster than $N^{-1/d}$ for any $d\geq 2$ (Theorem 1). Furthermore, in Section 3, we consider the Halton-type sequences in the sense of Hofer [Ho13, Ho18], and show that some of those sequences are also not quasi-uniform. These include the Faure sequence [Fa82], which was already disproven to be quasi-uniform in [DGS25]. Our purpose here is to provide an alternative proof from the Halton-type point of view.
@@ 1.1. Some known results
The covering radius has been studied under the name of dispersion [Nie92, Section 6]. It is known that certain classes of low-discrepancy constructions, such as $(t,d)$-sequences and the Halton sequence, achieve the optimal rate $N^{-1/d}$; see [Nie92, Theorems 6.11 & 6.12].
The separation radius, however, has not been studied to the same extent, and it has therefore remained unclear whether common low-discrepancy constructions are quasi-uniform. This question was explicitly raised in [WBG21]. In [SS07], Sobol’ and Shukhman studied the separation radius of the Sobol’ sequence [So67] and, based on numerical experiments, suggested that it also attains the optimal rate $N^{-1/d}$ for the separation radius. In the case $d=2$, however, this prediction was disproven recently in [G24, S25]. This shows that the Sobol’ sequence is not quasi-uniform, at least in two dimensions, whereas the question remains open for $d\geq 3$.
Further examples of low-discrepancy point sets and sequences that are not quasi-uniform can be found in [DGS25], including the Faure sequence. Examples of low-discrepancy point sets and sequences that are quasi-uniform are given in [DGPLS25, DGS25, G24b], in addition to the one-dimensional van der Corput sequence in any base $b$. The latter result means that any one-dimensional projection of the Halton sequence is quasi-uniform, whereas our main result below implies that no higher-dimensional projection satisfies this property.
@@ 2. Halton sequences are not quasi-uniform
Throughout this section, let $b_{1},\ldots,b_{d}$ be pairwise relatively prime integers with $1<b_{1}<\cdots<b_{d}$. We denote Euler’s totient of $b_{1}$ by $r$. Note that $b_{j}^{r}\equiv 1\pmod{b_{1}}$ holds for $2\leq j\leq d$. In this section, we prove the following result.
@@
Let $d\geq 2$, and let $b_{1},\dots,b_{d}$ be pairwise relatively prime integers. Let $r$ be Euler’s totient function of $b_{1}$. Let $\ell$, $k$, $c_{2},\dots,c_{d}\in\mathbb{N}$, and define
@@
For $d=2$ or $j=2$, the inequalities (1) hold trivially. Hence, we assume $d\geq 3$. Let $\alpha_{j}=\log b_{2}/\log b_{j}$ for $j=3,\dots,d$. By Dirichlet’s simultaneous approximation theorem [Sch80, Chapter 2], for any $k\in\mathbb{N}$, there exist integers $1\leq c_{2}\leq b_{1}^{k(d-2)}$ and $c_{j}\in\mathbb{Z}$ for $3\leq j\leq d$ such that
@@
For $k\in\mathbb{N}$, let $(c_{2},\dots,c_{d}):=(c_{2}(k),\dots,c_{d}(k))$ be integers satisfying (1) with $c_{j}\leq b_{1}^{k(d-2)}$, and define
@@
First, we show that $0\leq n<m<N$ holds. Using Proposition 2, we obtain
@@
Next, we derive an inequality needed for the proof. Assume that $k$ is large enough such that $b_{1}^{k}\geq 2(b_{1}+2)b_{d}^{dr}$. Then we have
@@
Finally, we bound the distance between the points $\boldsymbol{x}_{n}$ and $\boldsymbol{x}_{m}$. From Proposition 1 and (1), for $2\leq j\leq d$, we have
$|\varphi_{b_{j}}(n)-\varphi_{b_{j}}(m)|\leq\frac{1}{b_{j}^{rc_{j}b_{1}^{k}}}\leq\frac{b_{d}^{r}}{b_{2}^{rc_{2}b_{1}^{k}}}=\dfrac{b_{d}^{2r}}{M}$
and
$|\varphi_{b_{1}}(n)-\varphi_{b_{1}}(m)|\leq\frac{2}{b_{1}^{\ell+k+1}}\leq\frac{2}{M}.$
Hence, using (3), under the assumption $b_{1}^{k}\geq 2(b_{1}+2)b_{d}^{dr}$, we have
@@
It remains an open question whether this upper bound can be improved, i.e., whether a sharper estimate on the separation radius is possible.
@@ 3. Further results: Halton-type sequences
As mentioned in Section 1, the Faure sequence is known not to be quasi-uniform. More precisely, for any prime base $p$, the $p$-dimensional Faure sequence fails to be quasi-uniform; see [DGS25, Theorem 4.3]. The proof in [DGS25] relies on properties of Pascal matrices modulo $p$ together with Lucas’s theorem. In this section, we work out an alternative proof of [DGS25, Theorem 4.3] from the viewpoint of Halton-type sequences, using arithmetic in the ring of polynomials. We also present further examples of Halton-type sequences that are not quasi-uniform.
@@ 3.1. Definitions
To begin with, we introduce a polynomial analogue of the radical inverse function $\varphi_{b}$ and recall the definition of Halton-type sequences in the sense of Hofer [Ho13]. For more general definitions of Halton-type sequences, we refer to [Ho18]. In what follows, $\mathbb{P}$ denotes the set of prime numbers. For $p\in\mathbb{P}$, let $\mathbb{F}_{p}$ be the finite field with $p$ elements, identified with the set $\{0,1,\ldots,p-1\}$ equipped with addition and multiplication modulo $p$. Moreover, we write $\mathbb{F}_{p}[X]$ for the ring of polynomials over $\mathbb{F}_{p}$.
@@
Let $p\in\mathbb{P}$, and let $b(X)\in\mathbb{F}_{p}[X]$ be a polynomial of degree $\deg(b(X))=:e\geq 1$. For $n\in\mathbb{N}_{0}$ with its $p$-adic expansion $n=n_{0}+n_{1}p+n_{2}p^{2}+\cdots$, we associate the polynomial $n(X)=n_{0}X^{0}+n_{1}X^{1}+n_{2}X^{2}+\cdots\in\mathbb{F}_{p}[X]$. We then expand $n(X)$ in base $b(X)$, that is
@@
Additionally, we introduce the digital construction and $(t,d)$-sequences in a reduced form, which is sufficient for our purposes (for more details, see [Nie92, Chapter 4]).
@@
Define the Pascal matrix $P$ as $P=(\binom{j}{i})_{i,j\geq 0}$, where we set $\binom{j}{i}=0$ if $j<i$. Let $I$ be the $\infty\times\infty$ identity matrix. For a prime $p\in\mathbb{P}$, set $C_{1}:=I$, $C_{2}:=P\pmod{p}$, $\ldots$, $C_{p}:=P^{p-1}\pmod{p}$. Then the Faure sequence in base $p$ is the digital sequence generated by $C_{1},\ldots,C_{p}$, which is also a $(0,p)$-sequence over $\mathbb{F}_{p}$.
Similarly, let $b_{1}(X)=X$, $b_{2}(X)=X-1$, $\ldots$, $b_{p}(X)=X-(p-1)$ over $\mathbb{F}_{p}$. Then the Faure sequence in base $p$ is the Halton-type sequence in these bases (see, e.g., [Ho13, Example 4]). Moreover, in the case $p=2$, the corresponding Halton-type sequence coincides with the two-dimensional Sobol’ sequence.
Before moving on, we point out that, as proven in [Ho18, Proof of Theorem 4], any Halton-type sequence is a $(0,\boldsymbol{e},d)$-sequence in base $p$ with $\boldsymbol{e}=(e_{1},\ldots,e_{d})$ in the sense of Tezuka [Te13], who introduced the concept of $(t,\boldsymbol{e},d)$-sequences.
@@ 3.2. Results
The argument follows the proof of [Nie92, Theorem 6.12]. For each $\ell\in\{1,\ldots,d\}$, let $j_{\ell}$ be the unique non-negative integer such that $p^{e_{\ell}j_{\ell}}\leq N^{1/d}<p^{e_{\ell}(j_{\ell}+1)}$. Consider the partition of $0,1)^{d}$ into intervals of the form
@@
with integers $0\leq c_{\ell}<p^{e_{\ell}j_{\ell}}$ for $1\leq\ell\leq d$. As shown in [Ho18, Proof of Theorem 4], among the first $N^{\prime}=\prod_{\ell=1}^{d}p^{e_{\ell}j_{\ell}}$ points of $\mathcal{S}$, exactly one point lies in each such interval $J$. Hence
@@
The first case shows that any two-dimensional projection of the successive coordinates of the Faure sequence is not quasi-uniform. In particular, when $p=2$, the cases $a=0$ and $a=1$ coincide, and the corresponding sequence is the two-dimensional Sobol’ sequence that is already known to be not quasi-uniform [G24, S25]. Thus, this result can be seen as a generalization of the known result.
The second case is related to the three-dimensional Sobol’ sequence, as both sequences are $(0,(1,1,2),3)$-sequences constructed from the same polynomials $X,X+1,X^{2}+X+1$ over $\mathbb{F}_{2}$. Hence, the result can be seen as an extension of [G24, S25].
The last case corresponds exactly to the Faure sequence. Since this sequence is already known to be not quasi-uniform [DGS25], our contribution here is to provide an alternative proof from the viewpoint of Halton-type sequences.
@@
The proof of the third case relies on the equality $d=p$; that is, our argument works only when all components of the Faure sequence in base $p$ are considered. This is also the case in [DGS25].
Furthermore, our bound on $|\varphi_{b_{\ell}(X)}(n_{1})-\varphi_{b_{\ell}(X)}(n_{2})|$ for $\ell\in\{2,\ldots,p\}$ relies on the fact that $\boldsymbol{x}_{n_{1}}$ and $\boldsymbol{x}_{n_{2}}$ share the first $p^{w}-1$ or $p^{w}$ digits in their base $p$ expansions. From our proof, we can easily derive a generalization of [DGS25, Theorem 4.2]: let $L_{2},L_{3},\ldots,L_{p-1}$ be any non-singular lower triangular matrices over $\mathbb{F}_{p}$, then the $(0,p)$-sequence constructed by $(I,L_{2}P,\ldots,L_{p}P^{p-1})\pmod{p}$ satisfies

### v2

# Disproving the quasi-uniformity of the Halton sequences and of some Halton-type sequences

@@ Abstract.
In this short article, we prove that the Halton sequence, one of the most well-known low-discrepancy sequences, is not quasi-uniform in any dimension $d\geq 2$ with any pairwise relatively prime bases. We further disprove the quasi-uniformity of some Halton-type sequences, including the $p$-dimensional Faure sequence in base $p$, $p\in\mathbb{P}$, which provides an alternative proof of the known results.
@@ 1. Introduction
The Halton sequence is evidently one of the most well-known low-discrepancy sequences in the unit cube $0,1]^{d}$ for any dimension $d\geq 1$ [15, 19]. Throughout this paper, for an integer $b\geq 2$, we abbreviate the base-$b$ representation of $n=n_{0}+n_{1}b+n_{2}b^{2}+\cdots\in\mathbb{N}_{0}$ as $(n_{0},n_{1},n_{2},\ldots)_{b}$, where $\mathbb{N}_{0}$ denotes the set of non-negative integers and $n_{0},n_{1},\ldots\in\{0,1,\ldots,b-1\}$. The radical inverse function $\varphi_{b}$ in base $b$ is then defined by
@@
The discrepancy of the Halton sequence was studied already by Halton himself in [9], who proved an upper bound of $\mathcal{O}(N^{-1}(\log N)^{d})$ for the star discrepancy of the initial $N$ points of the sequence. Much later, Atanassov [1] provided a refined analysis, which led to a substantial improvement of the implied constant in this $\mathcal{O}$-bound. While Halton’s original constant grows super-exponentially fast in the dimension $d$, Atanassov’s constant decreases to 0 super-exponentially. In [16, 17], Levin proved a matching-order lower bound for the star discrepancy of the Halton sequence. Moreover, the tractability property of the weighted star discrepancy of Halton sequences has also been studied in [11].
With these results, together with the classical Koksma–Hlawka inequality [14, Chapter 2], the use of Halton sequences in high-dimensional numerical integration is theoretically well justified. However, Halton sequences have also been employed as sampling nodes for scattered data approximation—particularly in the context of radial basis function approximation (see, for instance, [5, 18, 23]). In such applications, not the low discrepancy but rather the geometric properties of the point set, such as the coverage of the underlying domain and the spacing between points, are crucial, as they contribute to bounds on the approximation error and to numerical stability [2, 21, 29, 30]. These geometric properties can be quantified in terms of the covering radius and the separation radius of the point set, naturally leading to the notion of quasi-uniformity.
@@
see, for instance, [3, Appendix A.1]. Thus, the quasi-uniformity requires that both radii decay at the optimal rate of order $N^{-1/d}$. As shown, for instance, in [19, Theorem 6.12], it is known that for $\mathcal{S}$ being the Halton sequence in any relatively prime bases $b_{1},\ldots,b_{d}\geq 2$,
@@
In passing, it is known that the first point $\boldsymbol{x}_{0}$, which lies at the origin, is well separated from the subsequent points of the Halton sequence [10, 20]. Let $x_{i,\ell}$ denote the $\ell$-th coordinate of the $i$-th point $\boldsymbol{x}_{i}$. In the present context, this separation property can be rephrased as the existence of a constant $c_{d}>0$, depending only on the bases $b_{1},\ldots,b_{d}$, such that
@@
Here, the first inequality follows from the arithmetic–geometric mean inequality, and the second inequality is from [20, Theorem 3.1].
One of the main purposes of this article is to disprove the quasi-uniformity of the Halton sequences in any dimension $d\geq 2$. This is done in Section 2 by proving that the separation radius of the Halton sequence indeed decays faster than $N^{-1/d}$ for any $d\geq 2$ (Theorem 1). Furthermore, in Section 3, we consider the Halton-type sequences in the sense of Hofer [12, 13], and show that some of those sequences are also not quasi-uniform. These include the $p$-dimensional Faure sequence in base $p$ with $p\in\mathbb{P}$ [6], which was already disproven to be quasi-uniform in [4]. Our purpose here is to provide an alternative proof from the Halton-type point of view.
@@ 1.1. Some known results
The covering radius has been studied under the name of dispersion [19, Section 6]. It is known that certain classes of low-discrepancy constructions, such as $(t,d)$-sequences and the Halton sequence, achieve the optimal rate $N^{-1/d}$; see [19, Theorems 6.11 & 6.12].
The separation radius, however, has not been studied to the same extent, and it has therefore remained unclear whether common low-discrepancy constructions are quasi-uniform. This question was explicitly raised in [31]. In [24], Sobol’ and Shukhman studied the separation radius of the Sobol’ sequence [25] and, based on numerical experiments, suggested that it also attains the optimal rate $N^{-1/d}$ for the separation radius. In the case $d=2$, however, this prediction was disproven recently in [8, 26]. This shows that the Sobol’ sequence is not quasi-uniform, at least in two dimensions, whereas the question remains open for $d\geq 3$.
Further examples of low-discrepancy point sets and sequences that are not quasi-uniform can be found in [4], including the $p$-dimensional Faure sequence in base $p$ with $p\in\mathbb{P}$. Examples of low-discrepancy point sets and sequences that are quasi-uniform are given in [3, 4, 7], in addition to the one-dimensional van der Corput sequence in any base $b$. The latter result means that any one-dimensional projection of the Halton sequence is quasi-uniform, whereas our main result below implies that no higher-dimensional projection satisfies this property.
@@ 2. Halton sequences are not quasi-uniform
Throughout this section, let $b_{1},\ldots,b_{d}$ be pairwise relatively prime integers with $1<b_{1}<\cdots<b_{d}$. We denote Euler’s totient function of $b_{1}$ by $r$, i.e. $r=\phi(b_{1})$. Note that then $b_{j}^{r}\equiv 1\pmod{b_{1}}$ holds for $2\leq j\leq d$ by Euler’s Theorem. In this section, we prove the following result.
@@
Let $d\geq 2$, and let $b_{1},\dots,b_{d}$ be pairwise relatively prime integers. We abbreviate Euler’s totient function of $b_{1}$ by $r$, i.e., $\phi(b_{1})=:r$. Let $\ell$, $k$, $c_{2},\dots,c_{d}\in\mathbb{N}$, and define
@@
For $d=2$ or $j=2$, the inequalities (1) hold trivially. Hence, we assume $d\geq 3$. Let $\alpha_{j}=\log b_{2}/\log b_{j}$ for $j=3,\dots,d$. By Dirichlet’s simultaneous approximation theorem [22, Chapter 2], for any $k\in\mathbb{N}$, there exist integers $1\leq c_{2}\leq b_{1}^{k(d-2)}$ and $c_{j}\in\mathbb{Z}$ for $3\leq j\leq d$ such that
@@
We abbreviate Euler’s totient function of $b_{1}$ by $r$, i.e. $\phi(b_{1})=:r$. For any $k\in\mathbb{N}$, let $(c_{2},\dots,c_{d}):=(c_{2}(k),\dots,c_{d}(k))$ be integers satisfying (1) with $c_{j}\leq b_{1}^{k(d-2)}$, and define
@@
First, we show that $0\leq n<m<N$ holds. Using Proposition 2 together with the definitions of $n$, $M$, $\ell$, and $N$, we obtain the following chain of inequalities:
@@
Next, we derive the inequalities in (3) that are needed for the proof. Choose $k$ sufficiently large so that $b_{1}^{k}\geq 2(b_{1}+2)b_{d}^{rd}$. Then we have
@@
Finally, we bound the distance between the points $\boldsymbol{x}_{n}$ and $\boldsymbol{x}_{m}$. From Proposition 1, the definition of $M$, and (1), for $2\leq j\leq d$, we derive
$|\varphi_{b_{j}}(n)-\varphi_{b_{j}}(m)|\leq\frac{1}{b_{j}^{rc_{j}b_{1}^{k}}}\leq\frac{b_{d}^{r}}{b_{2}^{rc_{2}b_{1}^{k}}}=\dfrac{b_{d}^{2r}}{M}.$
Proposition 1, together with the definition of $\ell$, implies that
$|\varphi_{b_{1}}(n)-\varphi_{b_{1}}(m)|\leq\frac{2}{b_{1}^{\ell+k+1}}\leq\frac{2}{b_{1}^{\log_{b_{1}}(M/b_{1}^{k})+k}}=\frac{2}{M}.$
Hence, using (3), which holds true under the assumption that $k$ is sufficiently large so that $b_{1}^{k}\geq 2(b_{1}+2)b_{d}^{dr}$, we have
@@
It remains an open question to determine the exact order of the separation radius for the Halton sequence.
@@ 3. Further results: Halton-type sequences
As mentioned in Section 1, for any prime base $p$, the $p$-dimensional Faure sequence fails to be quasi-uniform; see [4, Theorem 4.3]. The proof in [4] relies on properties of Pascal matrices modulo $p$ together with Lucas’s theorem. In this section, we work out an alternative proof of [4, Theorem 4.3] from the viewpoint of Halton-type sequences, using arithmetic in the ring of polynomials. We also present further examples of Halton-type sequences that are not quasi-uniform.
@@ 3.1. Definitions
To begin with, we introduce a polynomial analogue of the radical inverse function $\varphi_{b}$ and recall the definition of Halton-type sequences in the sense of Hofer [12]. For more general definitions of Halton-type sequences, we refer to [13]. Note that the notion of Halton-type sequences is different from that of polynomial Halton sequences defined by Tezuka [27]. In what follows, $\mathbb{P}$ denotes the set of prime numbers. For $p\in\mathbb{P}$, let $\mathbb{F}_{p}$ be the finite field with $p$ elements, identified with the set $\{0,1,\ldots,p-1\}$ equipped with addition and multiplication modulo $p$. Moreover, we write $\mathbb{F}_{p}[X]$ for the ring of polynomials over $\mathbb{F}_{p}$.
@@
Let $p\in\mathbb{P}$, and let $b(X)\in\mathbb{F}_{p}[X]$ be a polynomial of degree $\deg(b(X))=:e\geq 1$, $e\in\mathbb{N}$. For $n\in\mathbb{N}_{0}$ with its $p$-adic expansion $n=n_{0}+n_{1}p+n_{2}p^{2}+\cdots$, we associate the polynomial $n(X)=n_{0}X^{0}+n_{1}X^{1}+n_{2}X^{2}+\cdots\in\mathbb{F}_{p}[X]$. We then expand $n(X)$ in base $b(X)$, that is
@@
Additionally, we introduce the digital construction and $(t,d)$-sequences in a reduced form, which is sufficient for our purposes (for more details, see [19, Chapter 4]).
@@
Define the Pascal matrix $P$ as $P=(\binom{j}{i})_{i,j\geq 0}$, where we set $\binom{j}{i}=0$ if $j<i$. Let $I$ be the $\infty\times\infty$ identity matrix. For a prime $p\in\mathbb{P}$, set $C_{1}:=I$, $C_{2}:=P\pmod{p}$, $\ldots$, $C_{p}:=P^{p-1}\pmod{p}$. Then the $p$-dimensional Faure sequence in base $p$ is the digital sequence generated by $C_{1},\ldots,C_{p}$, which is also a $(0,p)$-sequence over $\mathbb{F}_{p}$.
Similarly, let $b_{1}(X)=X$, $b_{2}(X)=X-1$, $\ldots$, $b_{p}(X)=X-(p-1)$ over $\mathbb{F}_{p}$. Then the $p$-dimensional Faure sequence in base $p$ is the Halton-type sequence in these bases (see, e.g., [12, Example 4]). Moreover, in the case $p=2$, the corresponding Halton-type sequence coincides with the two-dimensional Sobol’ sequence.
Before moving on, we point out that, as proven in [13, Proof of Theorem 4], any Halton-type sequence is a $(0,\boldsymbol{e},d)$-sequence in base $p$ with $\boldsymbol{e}=(e_{1},\ldots,e_{d})$ in the sense of Tezuka [28], who introduced the concept of $(t,\boldsymbol{e},d)$-sequences.
@@ 3.2. Results
The argument follows the proof of [19, Theorem 6.12]. For each $\ell\in\{1,\ldots,d\}$, let $j_{\ell}$ be the unique non-negative integer such that $p^{e_{\ell}j_{\ell}}\leq N^{1/d}<p^{e_{\ell}(j_{\ell}+1)}$. Consider the partition of $0,1)^{d}$ into intervals of the form
@@
with integers $0\leq c_{\ell}<p^{e_{\ell}j_{\ell}}$ for $1\leq\ell\leq d$. As shown in [13, Proof of Theorem 4], among the first $N^{\prime}=\prod_{\ell=1}^{d}p^{e_{\ell}j_{\ell}}$ points of $\mathcal{S}$, exactly one point lies in each such interval $J$. Hence
@@
The first case shows that any two-dimensional projection of the successive coordinates of the Faure sequence is not quasi-uniform. In particular, when $p=2$, the cases $a=0$ and $a=1$ coincide, and the corresponding sequence is the two-dimensional Sobol’ sequence that is already known to be not quasi-uniform [8, 26]. Thus, this result can be seen as a generalization of the known result.
The second case is related to the three-dimensional Sobol’ sequence, as both sequences are $(0,(1,1,2),3)$-sequences constructed from the same polynomials $X,X+1,X^{2}+X+1$ over $\mathbb{F}_{2}$. Hence, the result can be seen as an extension of [8, 26].
The last case corresponds exactly to the $p$-dimensional Faure sequence in base $p$. Since this sequence is already known to be not quasi-uniform [4], our contribution here is to provide an alternative proof from the viewpoint of Halton-type sequences.
@@
The proof of the third case relies on the equality $d=p$; that is, our argument works only when all components of the $p$-dimensional Faure sequence in base $p$ are considered. This is also the case in [4].
Furthermore, our bound on $|\varphi_{b_{\ell}(X)}(n_{1})-\varphi_{b_{\ell}(X)}(n_{2})|$ for $\ell\in\{2,\ldots,p\}$ relies on the fact that $\boldsymbol{x}_{n_{1}}$ and $\boldsymbol{x}_{n_{2}}$ share the first $p^{w}-1$ or $p^{w}$ digits in their base $p$ expansions. From our proof, we can easily derive a generalization of [4, Theorem 4.2]: let $L_{2},L_{3},\ldots,L_{p-1}$ be any non-singular lower triangular matrices over $\mathbb{F}_{p}$, then the $(0,p)$-sequence constructed by $(I,L_{2}P,\ldots,L_{p}P^{p-1})\pmod{p}$ satisfies